# Local map Construction Methods with SD map: A Novel Survey

### v1

@@ 4.1 Align
Due to the inherent errors in GPS signals and the influence of vehicle movement, both vectorized and rasterized SD map priors inevitably have spatial misalignment with the current BEV space, making it difficult to fully align the two. Therefore, before fusion, it is necessary to spatially align the SD map prior with the current BEV operating space. FlexMap[66] uses SLAM trajectory and corrected RTK trajectory to calculate offset and achieve spatial alignment. To solve this problem, PriorMap[36] sets up a KEA (Knowledge Embedding Alignment) module to embed SD map prior knowledge and align it with image features in space. Specifically, first, a feature extraction network is used to extract feature points from the image and feature points from SD map prior knowledge.
Figure 5: The Overview of P-mapnet
@@
Subsequently, these feature points are spatially matched using an alignment algorithm based on attention mechanism. Finally, the aligned feature points are further processed through a fusion transformer network. Enhanced the accuracy and robustness of Local map perception algorithms. Similarly, P-MapNet[33] first downsamples the rasterized SD map prior, and then introduces a multi head cross attention module that allows the network to use cross attention to determine the most suitable alignment position, effectively enhancing BEV features using SD map prior. As Fig. 5, P-MapNet’s ablation experiment shows that even in the case of weak alignment with BEV space, directly concatenating SD map prior information still improves the performance of the model. On this basis, adding CNN modules and multi head cross attention modules can further improve the performance of the model. This demonstrates the important role of SD map prior information in Local map perception tasks, even without strict alignment, simply adding rasterized SD map priors can improve model performance.
Figure 6: The Overview of LGmap
4.2 Fusion
@@
In order to align the features of different sensors, it is necessary to achieve fusion in BEV level features. The image BEV features are obtained from the look around image through the perspective conversion module. In SMERF, the SD map feature interacts with the BEV feature through cross attention mechanism. Firstly, the BEV feature is encoded into a query vector and initialized through self attention mechanism. Given the SD map of the scene, LGMap, As shown in Fig. 6 evenly sample along each of the polylines for a fixed number of points. With sinusoidal embedding, BEVFormer apply cross-attention between the SD map feature representation with features from vision inputs on each encoder layer. The SD map features are encoded into key and value vectors, which are then computed through Cross Attention to obtain the final fused camera, BEV features of SD map.
Figure 7: The fusion methods of BLOS-BEV
@@
In addition to the common fusion methods of attention mechanisms, BLOS-BEV shows in Fig. 7, explore different fusion schemes that combine visual BEV features and SD map semantics to achieve optimal representation and performance, three SD map fusion techniques were explored: addition, concatenation, and cross-attention. Although all fusion methods outperform those that do not use SD maps, cross attention fusion of SD maps performs the best on the nuScenes and Argorse datasets, demonstrating excellent generalization performance and outstanding performance over long distances (150-200m).
@@
In P-mapnet, point cloud information has been added, and the lidar point cloud has been voxelated and MLP processed to obtain the feature representation of each point, resulting in Lidar BEV. Fusion of Image BEV and Lidar BEV is used to obtain further fused BEV features. Further convolutional downsampling of the fused BEV features can alleviate the misalignment problem between image BEV features and LiDAR BEV features.
Figure 8: The Overview of MapEx
@@
Then, through cross attention mechanism, the SD map good feature interacts with the fused BEV feature, resulting in the final fusion of camera and lidar point cloud, BEV features of SD map. Similarly, MapVision and MapEX, as Fig. 9 and Fig. 8 used the SD map feature as the key and value, and the feature map formed from multi-perspective images as the query to perform cross-attention.
Figure 9: The Pipeline of MapVision
@@
To address these issues such as occlusion and limited sensing range can result in inaccuracies, RoadPainter present a novel SD map interaction module that effectively augments BEV features by incorporating beyond-visual-range in Fig. 10.
Figure 10: The fusion methods of RoadPainter
@@
EORN in Fig. 11 rasterize the SD map and generate a SD map in BEV. An SD encoder based on a ResNet-18 to extract SD map features. SD map feature is then interpolated and concatenated with the BEV feature from images BEV along the channel dimension. Fusion methods uses a simple two-layer convolutional neural network, the ConvFuser, which fuses concatenated features and outputs the fused BEV features. Another method involves a graph-based encoder that fuses SD map graphs with BEV features and combines these with the outputs from the centerline deformable decoder using a multi-head attention mechanism.
Figure 11: The fusion methods of ENOR
@@
The subsequent decoder can calculate and output corresponding results for different tasks through query queries from BEV features containing rich information.
5 Conclusion and Discussions
This section explores the challenges of Local maps construction with SD map and points out the future direction. Lastly, this research make conclusions.
5.1 Challenges and Future Prospects
@@
Inference of Road Topological Relationships. The topological relationships in Local maps can be divided into two branches: the topological relationships between roads (primarily representing road connectivity) and the topological relationships between roads and traffic signs (including traffic control signals and other directional signs). Enhancing scene understanding of the road environment is crucial for high-level autonomous driving tasks. The OpenLane-v2 dataset is the first public dataset providing topological relationships between roads and between roads and traffic signs. Current research focusing on this area is still limited. Future work could model the topological structures of road networks and the scene understanding tasks of traffic signs using graph neural network models.
Incorporating More SD map Prior Information. Existing research has demonstrated that incorporating more road type information can enhance model performance. However, beyond the basic road network positions and road types, SD map can provide richer prior information. For example, OpenStreetMap offers additional information such as the number of lanes, lane directions, and road topological relationships. Future research could attempt to integrate this diverse information as SD map priors to further enhance the robustness and accuracy of Local map perception models.
5.2 Conclusions
In this article, the literature on local map construction using SD maps was reviewed, highlighting the pivotal role of SD maps in this task. The definition and core aspects of local map construction with SD maps were presented, demonstrating their significance in developing accurate and reliable maps. Commonly used public datasets and their corresponding evaluation metrics were enumerated.
The main processes of leading technological approaches were summarized, focusing on the representation and encoding methods for data from various sensors, such as LiDAR, cameras, and radar. Advanced fusion techniques for integrating multi-source sensor data, fusion methods, were explored, along with their respective strengths and limitations.
The evaluation prospects and design trends of local map construction models were discussed. This included addressing emerging challenges, such as improving SD map alignment with BEV perspectives and enhancing encoding and processing methods. The potential of incorporating detailed SD map prior information to model road topological relationships was considered, with the goal of improving scene understanding and supporting higher-level autonomous driving tasks.

### v2

@@ 4.1 Alignment Methods
Due to the inherent errors in GPS signals and the influence of vehicle movement, both vectorized and rasterized SDMap priors inevitably have spatial misalignment with the current BEV space, making it difficult to fully align the two. Therefore, before fusion, it is necessary to spatially align the SDMap prior with the current BEV operating space. FlexMap[66] uses SLAM trajectory and corrected RTK trajectory to calculate the offset and achieve spatial alignment.
Figure 19: The alignment method of PriorLane.
To solve this problem, PriorMap[35] sets up a KEA module to embed SDMap prior knowledge and align it with image features in space shown in Fig. 19. Specifically, first, a feature extraction network is used to extract feature points from the image and feature points from the SDMap prior knowledge. Subsequently, these feature points are spatially matched using an alignment algorithm based on attention mechanisms.
@@
Finally, the aligned feature points are further processed through a fusion transformer network and enhance the accuracy and robustness of local map perception algorithms. Similarly, P-MapNet[31] first downsamples the rasterized SDMap prior and then introduces a multi-head cross-attention module that allows the network to use cross-attention to determine the most suitable alignment position, effectively enhancing BEV features using SDMap prior, As illustrated in Fig. 20.
Figure 20: The alignment and fusion methods of P-MapNet using cross attention.
P-MapNet’s ablation experiment[31] shows that even in the case of weak alignment with BEV space, directly concatenating SDMap prior information still improves the performance of the model. On this basis, adding CNN modules and multi-head cross-attention modules can further improve the performance of the model. This demonstrates the important role of SDMap prior information in local map perception tasks, even without strict alignment, simply adding rasterized SDMap priors can improve model performance.
In LSM[70], to achieve cross-modality alignment, an additional SDMap cross-attention layer is introduced after each lane attention layer. This layer incorporates SDMap features, enhancing the model’s ability to align the spatial context between the SDMap and BEV features, leading to more accurate lane detection and topological reasoning.
4.2 Fusion Methods
@@
To align the features of different sensors, it is necessary to achieve fusion in BEV-level features. The image BEV features are obtained from the rounded image through the perspective conversion module. In SMERF[2], the SDMap feature interacts with the BEV feature through a cross-attention mechanism as demonstrated in Fig. 21.
Figure 21: The fusion method of SMERF.
@@
Firstly, the BEV feature is encoded into a query vector and initialized through a self-attention mechanism. Given the SDMap of the scene, LGMap[69] evenly samples along each of the polylines for a fixed number of points. With sinusoidal embedding, BEVFormer[18] applies cross-attention between the SDMap feature representation with features from vision inputs on each encoder layer. The SDMap features are encoded into key and value vectors, which are then computed through cross-attention to obtain the final fused camera, BEV features of SDMap.
@@
In addition to the common fusion method of attention mechanisms, BLOS-BEV[65] shows in Fig. 22, explore different fusion schemes that combine visual BEV features and SDMap semantics to achieve optimal representation and performance, three SDMap fusion techniques were explored: addition, concatenation, and cross-attention. Although all fusion methods outperform those that do not use SDMap, the cross-attention fusion of SDMap performs the best on the nuScenes[10] and Argoverse datasets[12], demonstrating excellent generalization performance and outstanding performance over long distances (150-200m).
Figure 22: The three kinds of fusion methods in BLOS-BEV.
@@
In P-MapNet[31], point cloud information has been added, and the lidar point cloud has been voxelated and MLP processed to obtain the feature representation of each point, resulting in lidar BEV. Fusion of Image BEV and lidar BEV are used to obtain further fused BEV features. Further convolutional downsampling of the fused BEV features can alleviate the misalignment problem between image BEV features and lidar BEV features.
Figure 23: The fusion method of MapVision using attention modules.
@@
Then, through cross attention mechanism, the SDMap feature interacts with the fused BEV feature, resulting in the final fusion of images and point cloud, BEV features of SDMap. Similarly, MapVision[36] uses the SDMap feature as the key and value, and the feature map formed from multi-perspective images as the query to perform cross-attention (Fig. 23).
Similar to LSM[70], a dedicated SDMap encoder is developed to encode vectorized map elements, enriched by BEV features to provide enhanced spatial context. This encoding helps incorporate crucial road layout information and facilitates a more precise understanding of lane detection and traffic elements in the scene.
Figure 24: The fusion method of LGmap: Streaming-stacking strategy.
Fig. 24 shows the fusion method of LGmap[69]. The stacking strategy and streaming strategy are the same as StreamMapNet’s summary[7]. To demonstrate the effectiveness of long-range stacking for the streaming-stacking strategy in the figure, the stacking previous frame interval parameter is set to 2. The stacking strategy only fuses one previous frame in this figure. It may fuse more than one frame.
@@
To address these issues such as occlusion and limited sensing range can result in inaccuracies, RoadPainter[72] presents a novel SDMap interaction module that effectively augments BEV features by incorporating beyond-visual-range in Fig. 25.
Figure 25: The fusion method of RoadPainter using BEV fusion module.
Figure 26: The fusion method of EORN using Convfuser module.
@@
EORN[73] rasterizes and generates SDMap in BEV. An SD encoder based on a ResNet-18[20] to extract SDMap features. The SDMap feature is then interpolated and concatenated with the BEV feature from images BEV along the channel dimension. fusion method uses a simple two-layer convolutional neural network, the ConvFuser, which fuses concatenated features and outputs the fused BEV features. Another method involves a graph-based encoder that fuses SDMap graphs with BEV features and combines these with the outputs from the centerline deformable decoder using a multi-head attention mechanism.
@@
Fig. 26 shows the fusion method of EORN[73]. The subsequent decoder can calculate and output corresponding results for different tasks through query queries from BEV features containing rich information.
5 Challenges and Future Prospects
Despite significant progress in local map construction with SDMap, several challenges remain that hinder the full realization of its potential. Addressing these issues will require further innovation, interdisciplinary collaboration, and the development of new methodologies.
  1. 1)
@@ 5 Challenges and Future Prospects
  3. 3)
    Inference of Road Topological Relationships. The topological relationships in local map can be divided into two branches: the topological relationships between roads (primarily representing road connectivity) and the topological relationships between roads and traffic signs (including traffic control signals and other directional signs). Enhancing scene understanding of the road environment is crucial for high-level intelligent driving tasks. The OpenLane-v2 dataset[17] is the first public dataset providing topological relationships between roads and between roads and traffic signs. Current research focusing on this area is still limited. Future work could model the topological structures of road networks and the scene understanding tasks of traffic signs using graph neural network models.
  4. 4)
    Incorporating More SDMap Prior Information. Existing research has demonstrated that incorporating more road-type information can enhance model performance. However, beyond the basic road network positions and road types, an SDMap can provide richer prior information. For example, OpenStreetMap[8] offers additional information such as the number of lanes, lane directions, and road topological relationships. Future research could attempt to integrate this diverse information as SDMap priors to enhance the robustness and accuracy of local map perception models.
6 Conclusion
In this article, the literature on local map construction using SDMap was reviewed, highlighting the pivotal role of SDMap in this task. The definition and core aspects of local map construction with SDMap were presented, demonstrating their significance in developing accurate and reliable maps. Commonly used public datasets and their corresponding evaluation metrics were enumerated.
This study reviews the core processes of state-of-the-art technologies, concentrating on data representation and encoding methods for sensor data, including lidar, camera, and radar inputs. It also examines sophisticated sensor fusion strategies for consolidating data from diverse sensors, evaluating their comparative advantages and constraints.
The evaluation prospects and design trends of local map construction models were discussed. This included addressing emerging challenges, such as improving SDMap alignment with BEV perspectives and enhancing encoding and processing methods. The potential of incorporating detailed SDMap prior information to model road topological relationships was considered, with the goal of improving scene understanding and supporting higher-level intelligent driving tasks.